OpenStudy (anonymous):

How much metabolic energy is required for a 68kg person to run at a speed of 15km/hr for 15min ?

OpenStudy (anonymous):

I know that running at 15km/hr takes 1150W of Metabolic Power but I'm not sure where to go from there.

OpenStudy (anonymous):

Because watt is joule per second should it be as simple as the product of metabolic power and the time 15 minutes?

OpenStudy (anonymous):

E=W (+Q) W= FD e=1/2 mv^2 These are the equations that I think are relevant but I can't seem to "plug and chug" the numbers appropriately.

OpenStudy (anonymous):

I don't think so. I know that the answer has to be in kilojoules.

OpenStudy (anonymous):

Its all I can help with but the answer will be in kilojoules though

OpenStudy (anonymous):

how did you calculate your power

OpenStudy (anonymous):

Haha it's alright. I appreciate it. Thanks!

OpenStudy (anonymous):

The power was given to be based on a table in my book.

OpenStudy (anonymous):

Do you have an answer in the back of the book

OpenStudy (anonymous):

Unfortunately not, I already checked haha. It's an even number problem >.<

OpenStudy (anonymous):

Why do you think that it wouldn't be power times time because that would give you 1350 j

OpenStudy (anonymous):

I already tried inputting 1350J as the answer but it was marked as incorrect.

OpenStudy (anonymous):

I also tried inputting 1.35kj and that was incorrect as well.

OpenStudy (anonymous):

Well no worries, I'll just ask my professor tomorrow. Thanks for the help!

OpenStudy (anonymous):

How did you get it to 1.35kj? By my calculations it should be 1035kJ

OpenStudy (anonymous):

whoops I meant 1350 kj

OpenStudy (anonymous):

I dropped the k

OpenStudy (anonymous):

t = 15 = 900s \[E _{k} = \frac{ (\frac{ 1 }{ 2 } \times m \times v ^{2} \times t) }{ t }\]\[E _{k} = \frac{ (\frac{ 1 }{ 2 } \times 68 \times 15 ^{2} \times 900) }{ 900 }\] \[E _{k} = 7650J\] There should be your answer :)